OpenStudy (anonymous):

(b) A certain computer simulation generates a random amount of data whenever it runs. On average it creates 30Mb of data, with a standard deviation of 5Mb. The simulation is run 64 times. Let D be the random variable for the total amount of data generated. (i) Find E(D) (ii) Find Var(D) (iii) Using a normal approximation, estimate the probability that the data will overflow the computer’s 2000Mb storage disc.

OpenStudy (anonymous):

(i) E(D)=nu=64*30=1920 (ii)Var(D)=n σ^2=64*5^2=64*25=1600 I don't know how to find part (iii)